# Kolmogorov-Arnold Network

### v2

@@ V-D Ablation Study
TABLE VI: Results of training a $\mathbb{C}$KAN (with one intermediate layer of size 1) on the full knot dataset ($15$ features), and using only the most relevant $7$ or $3$ features or leaving these out.
@@
TABLE VII: Ablation study on the influence of normalization scheme and $\mathbb{C}$SiLU type. Evaluated on a $15{\times}1{\times}14$ $\mathbb{C}$KAN on the knot dataset.
@@ VI Discussion
We have shown that $\mathbb{C}$KAN has a clear advantage over the real-valued KANs when dealing with complex-valued function-fitting tasks. Our approach can compete with KAN and FastKAN while having less parameters and a shallower network architecture. These properties affect the explainability in a positive way. Furthermore our $\mathbb{C}$KAN has proven to be more stable w.r.t. outliers in regression tasks since $\mathbb{C}$KAN always produces one of the best MSE scores but is sometimes outperformed on the MAE metric. Our approach also produces more stable results across the five different runs of cross-validation for each configuration, which underlines the consistently good generalization capabilities of our model.
While $\mathbb{C}$KAN does not excel on datasets with mostly real-valued features, like the circuit dataset in subsection V-B, it still performs similarly well as the real-valued KANs. If the datasets contain more complex numbers, $\mathbb{C}$KAN can unfold it’s full potential and outperform the real-valued KANs while requiring less parameters and having a shallower network structure, which improves the explainability.
Our tool for visualizing the proposed $\mathbb{C}$KAN enables the understanding of the calculations happening inside our $\mathbb{C}$KAN and can thus help with (re-) discovering mathematical correlations in the data like Liu et al. pointed out in their real-valued KAN [1, 2]. The calculated importance scores reflect the true relevances of the features, as is shown by our experiments of training only on the $3$ or $7$ most important features or by leaving them out in Table VI. While the plot of our $\mathbb{C}$KAN trained on synthetic formula $f_{4}$ in Fig. 3 may seem unintuitive at first glance, the learned functions that are plotted show an interesting insight. In the bottom layer there exist two roots in both functions for $z_{1}^{2}$ and $z_{2}^{2}$. The root points are exactly rotated around $90$° when comparing those two functions. This is caused by shifts that cancel out:
@@ VII Conclusion
In this work we have developed $\mathbb{C}$KAN, a complex-valued variant of KAN, which relies on complex-valued RBFs. We have introduced BatchNorm to efficiently meet the problem of outputs exceeding the grid size of subsequent layers. In our experiment we could show that $\mathbb{C}$KAN outperforms the real-valued version KAN and FastKAN consistently on complex-valued tasks and performs on par or superior on tasks that mix complex- and real-valued inputs. Additionally, $\mathbb{C}$KAN training is more stable and leads to more explainable models, since less layers are needed to produce good results.
As future work, we aim to apply $\mathbb{C}$KAN to real world problems with complex-valued data, such as finding solutions to complex-valued PDEs by utilizing them in Kolmogorov-Arnold-Informed neural networks [4].

### v3

@@ V-D Ablation Study
TABLE VI: Results of training a CVKAN (with one intermediate layer of size 1) on the full knot dataset ($15$ features), and using only the most relevant $7$ or $3$ features or leaving these out.
@@
TABLE VII: Ablation study on the influence of normalization scheme and $\mathbb{C}$SiLU type. Evaluated on a $15{\times}1{\times}14$ CVKAN on the knot dataset.
@@ VI Discussion
We have shown that CVKAN has a clear advantage over the real-valued KANs when dealing with complex-valued function-fitting tasks. Our approach can compete with KAN and FastKAN while having less parameters and a shallower network architecture. These properties affect the explainability in a positive way. Furthermore our CVKAN has proven to be more stable w.r.t. outliers in regression tasks since CVKAN always produces one of the best MSE scores but is sometimes outperformed on the MAE metric. Our approach also produces more stable results across the five different runs of cross-validation for each configuration, which underlines the consistently good generalization capabilities of our model.
While CVKAN does not excel on datasets with mostly real-valued features, like the circuit dataset in subsection V-B, it still performs similarly well as the real-valued KANs. If the datasets contain more complex numbers, CVKAN can unfold it’s full potential and outperform the real-valued KANs while requiring less parameters and having a shallower network structure, which improves the explainability.
Our tool for visualizing the proposed CVKAN enables the understanding of the calculations happening inside our CVKAN and can thus help with (re-) discovering mathematical correlations in the data like Liu et al. pointed out in their real-valued KAN [1, 2]. The calculated importance scores reflect the true relevances of the features, as is shown by our experiments of training only on the $3$ or $7$ most important features or by leaving them out in Table VI. While the plot of our CVKAN trained on synthetic formula $f_{4}$ in Fig. 3 may seem unintuitive at first glance, the learned functions that are plotted show an interesting insight. In the bottom layer there exist two roots in both functions for $z_{1}^{2}$ and $z_{2}^{2}$. The root points are exactly rotated around $90$° when comparing those two functions. This is caused by shifts that cancel out:
@@ VII Conclusion
In this work we have developed CVKAN, a complex-valued variant of KAN, which relies on complex-valued RBFs. We have introduced BatchNorm to efficiently tackle the problem of outputs exceeding the grid size of subsequent layers. In our experiment we could show that CVKAN outperforms the real-valued versions KAN and FastKAN consistently on complex-valued tasks and performs on par or superior on tasks that mix complex- and real-valued inputs. Additionally, CVKAN training is more stable and leads to more explainable models, since less layers are needed to produce good results.
As future work, we aim to apply CVKAN to real world problems with complex-valued data, such as finding solutions to complex-valued PDEs by utilizing them in Kolmogorov-Arnold-Informed Neural Networks [4]. Additionally, the concept of CVKAN could be extended further to other arbitrary hypercomplex algebras such as quaternions [41, 42].